Two point charges +4q and +q are separated by distance r, where should a third point charge Q be placed so that the whole system remains in equilibrium?
1
Expert's answer
2020-04-13T09:58:10-0400

Let place the charge Q on the line between +4q and q charges as shown on the figure.





Then, according to the Coulomb's low, the interaction force between +4q and Q charges is:

"F_1 = k\dfrac{4q\cdot Q}{x^2}.\\"

The interaction force between +q and Q charges is:

"F_2 = k\dfrac{q\cdot Q}{(r - x)^2}, \\" where "r" is the distance between +4q and +q and "k" is the Coulomb's constant.

In the equilibrium state both forces are equal, thus obtain the equation:

"k\dfrac{4q\cdot Q}{x^2} = k\dfrac{q\cdot Q}{(r - x)^2}" .

Solving it for "x" , obtain:

"\dfrac{4}{x^2} = \dfrac{1}{(r - x)^2} \Rightarrow 4(r-x)^2 = x^2 \Rightarrow 2(r-x) = x \\
x = \dfrac{2}{3}r.\\"


Answer. The charge Q should be placed at the distance 2/3r from the charge +4q.